OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 + 24x + 62? x = –6 x = 6 x = –10 x = 10

myininaya (myininaya):

Use the formula x=-b/(2a) This will be axis of symmetry for a parabola in the form y=ax^2+bx+c.

myininaya (myininaya):

You to identify a and b.

myininaya (myininaya):

can you tell me what a is in your equation above. y=ax^2+bx+c y=2x^2+24x+c